OpenStudy (anonymous):

using cramer's rule solve this linear equation 2x+y+z=3 x-y-z=0 x+2y+z=0

OpenStudy (anonymous):

First step : formulating as Cramer's matrix : \[\begin{bmatrix} 2 & 1 & 1\\ 0 & -1 & -1\\ 1 & 2 & 1 \end{bmatrix} \begin{bmatrix} x\\ y\\ z \end{bmatrix} = \begin{bmatrix} 3\\ 0\\ 0 \end{bmatrix} \]

OpenStudy (anonymous):

then find the inverse of the matrix on the left... can you do that?

OpenStudy (anonymous):

lionida?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=sovle {{2%2C1%2C1}%2C{0%2C-1%2C-1}%2C{1%2C2%2C1}}*{{x}%2C{y}%2C{z}}%3D{{3}%2C{0}%2C{0}}